OpenStudy (anonymous):

y=2/3x-4 I have to graph this equation... after it is solved

hero (hero):

What do you mean after it is solved?

OpenStudy (anonymous):

i am finding the x and y intercept and then graphing it

hero (hero):

Ah, I see. That explains it..

hero (hero):

Well y = -4 x = 6

OpenStudy (anonymous):

how do you get the 6? are you multiplying 2by 3?

hero (hero):

No: 0 = 2/3x - 4 4 = 2/3x 12 = 2/x 6 = 1/x x = 1/6 my bad, it's not 6 , it's 1/6

OpenStudy (anonymous):

how do I graph 1/6?

hero (hero):



hero (hero):

Should look something like that

hero (hero):

You have to set the x scale.... let the first line after 0 on the x scale = 1/6, then 2/6, 5hen 3/6, 4/6, 5/6 and 6/6 = 1 let the y scale = -1, -2, -3, -4